# RobustSpring: Benchmarking Robustness to Image Corruptions for Optical Flow, Scene Flow and Stereo

### v1

@@ 4 Results
We evaluate RobustSpring under two aspects: First, we report initial results for 16 optical flow, scene flow and stereo models. Then, we analyze the benchmark evaluation, particularly subsampling strategy and ranking methods.
@@ 4.1 Initial RobustSpring Benchmark Results
We provide initial results on RobustSpring for selected models from all three dense matching tasks. For optical flow, we include GMFlow [76], MS-RAFT+ [27, 26], FlowFormer [23], GMA [28], SPyNet [52], RAFT [69], FlowNet2 [24], and PWCNet [65]. For scene flow, we evaluate M-FUSE [42] and RAFT-3D [70]. For stereo estimation, we evaluate RAFT-Stereo [36], ACVNet [75], LEAStereo [13], and GANet [79]. An overview of all models and used checkpoints is in Tab. A1. Importantly, none of these models are fine-tuned to either Spring or RobustSpring data, to assess the generalization capacity of existing models.
Optical Flow. The evaluation results in Tab. 2 show considerable robustness variations over the different corruption types, which we also visualize in Fig. 4(a). Weather-based corruptions, especially rain and snow, degrade the performance most and lead to the largest $R^{c}$ values. In contrast, color-based corruptions have a relatively small impact, as most models maintain low $R^{c}_{\text{EPE}}$ values. Also, the order of models can change significantly depending on the corruption type. While FlowNet2 does not perform well in the overall comparison, it is the best model for noise-based corruptions in Figure 4(b). Overall, GMFlow achieves the lowest average $R^{c}_{\text{EPE}}$, GMA the lowest median. We will detail on ranking differences in Sec. 4.2.
@@
Figure 4: Accumulated corruption robustness $R^{c}_{\text{EPE}}$ for optical flow models over all corruptions [top] and only noise corruptions [bottom]. All other corruption classes color (purple), blur (blue), noise (cyan), quality (green), and weather (yellow) are in Fig. A3.
Figure 5: Accuracy vs. robustness of optical flow models, measured as EPE and Averaged $R^{c}_{\text{EPE}}$. Small values indicate accurate and robust models. Fig. A2 shows accuracy vs. Median $R^{c}_{\text{EPE}}$.
@@
Table 3: Initial RobustSpring results on corruption robustness of scene flow models. Corresponding Disparity 1 from LEAStereo (s) for M-FUSE, and GANet (s) for RAFT-3D in Tab. 4.
To investigate a potential accuracy-robustness tradeoff on image corruptions, we visualize both quantities in 5. Overall, accurate models tend to be more robust, though we find a slight tradeoff because there is no unanimous winner in both dimensions – similarly for median robustness in Fig. A2. Interestingly, this contrasts with adversarial robustness evaluations, which observed a clear accuracy-robustness tradeoff on optical flow [60]. Potentially, this tradeoff is less pronounced for image corruptions as they are not optimized per model like adversarial attacks.
Focusing on the architecture of optical flow models, we find that transformer-based models, such as GMFlow and FlowFormer, generally outperform other architectures. However, they tend to struggle with noise corruptions, potentially resulting from their global matching. Hierarchical models, such as MS-RAFT+, achieve balanced performance for most corruptions and may benefit from multi-scale feature processing to cope with quality degradations. In contrast, stacked architectures such as FlowNet2 are uniquely resilient to noise, potentially due to their progressive refinement across layers. Overall, certain architectural features appear to influence robustness to certain corruption types, but there is no clear winner in terms of architecture.
@@
Table 4: Initial RobustSpring results on corruption robustness of stereo disparity models, using corruption robustness $R^{c}_{\text{1px}}$, $R^{c}_{\text{Abs}}$ and $R^{c}_{\text{Dl}}$ between clean and corrupted predictions. Low values indicate robust models, using Stereo (s) and KITTI (k) checkpoints, cf. Tab. A1.
Scene Flow. The results for scene flow are in Tab. 3, and include optical flow and target frame disparity predictions for M-FUSE and RAFT-3D. M-FUSE generally produces more robust optical flow across corruptions with a lower average $R^{c}_{\text{EPE}}$ than RAFT-3D. But both methods suffer significant performance losses for severe weather like rain and noise-based corruptions, e.g. impulse noise. Interestingly, their robustness does not improve compared to conventional optical flow models. Noise and weather corruptions remain a challenge for Disparity 2 predictions. Here, RAFT-3D consistently achieves lower robustness scores compared to M-FUSE, but conditions like impulse noise or rain still notably affect disparity predictions. Overall, both models have limited robustness, but temporal consistency may contribute to lower robustness scores under several corruption types.
Stereo. The results of the stereo disparity estimations are presented in Tab. 4. The effect of the different corruptions on the performance is significant, with noise and weather-based corruptions leading to the largest errors, especially for GANet and LEAStereo. In particular, Gaussian and impulse noise introduce extremely large errors, highlighting the sensitivity of stereo models to pixel-level noise. Blur distortions, especially zoom blur, also have a severe impact on all models, with high 1px and D1 errors. In contrast, color-based distortions generally yield smaller errors. RAFT-Stereo shows stronger resilience across most corruption groups, performing better on color and noise based corruption than other models. However, it also struggles with noise and severe weather effects such as rain and snow.
4.2 Metrics and Benchmark Capability
Subsampling. We evaluate RobustSpring’s strict data subsampling by comparing to results on the full test set. As shown in Tab. 5, our subsampling strategy produces results that are nearly identical to those that include all pixels in the robustness calculation. We observe the largest discrepancy for Spring’s original subsampling, because it includes a handful of full-resolution Hero-frames. If those frames are also subsampled (Spring*), results align with the full dataset. Overall, our stricter subsampling to 0.05% of all data is not only data efficient but also exact.
@@ 4.2 Metrics and Benchmark Capability
Table 5: Influence of subsampling. We compare robustness evaluations on the full test data (Full) to evaluations on Spring’s original subsampling (Spring), original subsampling without Hero-frames (Spring*), and our refined corruption subsampling (Ours).
Metric Ranking. To explore how ranking strategies influences the optical-flow robustness order, we contrast our three summarization strategies: Average, Median, and Schulze, cf. Sec. C.3. The rankings in Tab. 6 notably differ across strategies. The Average differs most from the other rankings. For example, it ranks GMFlow 1st, which is only 4th on Median and Schulze, suggesting a good performance across corruptions without excessive outliers but no top performance on most corruptions. Interestingly, Median and Schulze rankings are more aligned. As Schulze’s ranking involves complex comparisons of per-corruption rankings and must be globally recomputed for new models, the Median ranking is a cheap approximation to it. The ranking strategy has significant implications for selecting robust models. No model is optimal across rankings, and the rankings accentuate different aspects: overall performance, outlier robustness, or balanced performance in pairwise comparisons. Hence, RobustSpring reports them all.
@@
Table 6: Robustness ranking of optical flow models with ranking strategies Average $R^{c}_{\text{EPE}}$, Median $R^{c}_{\text{EPE}}$, and Schulze to summarize results over corruptions. Schulze does not produce numeric values.
Corruption Robustness on Object Corruptions.
@@
Figure 6: Stability of corruption robustness $R^{c}_{\text{EPE}}$ on rain corruption. Robustness scores and method rankings remain stable even if we exclude all rain pixels (90% of all pixels) from $R^{c}_{\text{EPE}}$ calculation, which places dominating prediction errors in the background beyond the moving particles. Results for snow are in Fig. A4.
Intuitively, models are robust if they recover the main scene despite image corruptions. Here, we investigate if the corruption robustness metric faithfully represents model robustness even if corruptions like rain introduce moving objects to the scene. To this end, we contrast the robustness score contributions of background and corruption objects, by excluding pixels of objects like rain drops from the score calculation. We detect object pixels by taking the value difference $d$ between original and corrupt images, and exclude them if $(1-d)$ is above a detection threshold. Threshold 0 detects no rain pixels, matching the vanilla $R^{\text{Rain}}_{\text{EPE}}$, while 100 detects all. Figure 6 shows the robustness score if rain is excluded from the calculation, along with bars indicating the amount [%] of excluded pixels. Remarkably, the robustness score is stable, i.e. varies $\leq$5%, even for discarding all rain pixels, i.e. 90% of all pixels. Large robustness scores on rain or snow, cf. Fig. A4, thus stem from mispredictions in the periphery of altered pixels, not from motion predictions on altered pixels. As it is dominated by scene-wide effects, our corruption robustness yields stable robustness rankings that make it suited for broad model robustness evaluations.
Robustness in the Real World.
@@
Figure 7: Relative robustness to noise on RobustSpring transfers to noisy real-world KITTI data [45] for most optical flow models.
@@
Finally, we investigate if RobustSpring’s corruption robustness transfers to the real world. To this end, we select the noisiest 10% KITTI data, estimating noise as in [25]. These noisy KITTI frames have no clean counterparts to calculate corruption robustness $R^{\text{Noise}}_{\text{EPE}}$. Thus, we approximate $R^{\text{Noise}}_{\text{EPE}}$ via the accuracy difference on noisy and non-noisy KITTI frames. To account for model-specific performance differences on Spring and KITTI, we normalize with the clean dataset performance and show the resulting relative robustness $\frac{R^{\text{Noise}}_{\text{EPE}}}{\text{EPE}^{\text{Clean}}}$ in Fig. 7. Relatively robust models with low scores on RobustSpring are also robust on KITTI and vice versa. The only outlier, FlowFormer, overperforms on KITTI, potentially due to outstanding memorization capacity and exposure to KITTI during training. Because overall noise resilience on RobustSpring qualitatively transfers to KITTI, RobustSpring supports model selection for real-world settings where corruption robustness cannot be measured.
@@ 5 Conclusion
With RobustSpring we introduce an image corruption dataset and benchmark that evaluates the robustness of optical flow, scene flow and stereo models. We carefully design 20 different image corruptions and integrate them in time, stereo, and depth for a holistic evaluation of dense matching tasks. Furthermore, we establish a corruption robustness metric using clean and corrupted predictions, and compare ranking strategies to unify model results across all 20 corruptions. RobustSpring’s benchmark further supports data-efficient result uploads to a public website. Our initial evaluation of 16 optical flow, scene flow and stereo models reveals an overall high sensitivity to corrupted images. As our robustness results translate to real-world performance, systematic corruption benchmarks like RobustSpring are crucial to uncover potential model performance improvements.
@@
Acknowledgements. Jenny Schmalfuss and Andrés Bruhn acknowledge support by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – Project-ID 533085500. Furthermore, Jenny Schmalfuss and Victor Oei acknowledge funding by the DFG – Project-ID 251654672 – TRR 161 (B04). Margret Keuper would like to acknowledge support by the DFG Research Unit 5336 – Learning to Sense (L2S). The authors thank the International Max Planck Research School for Intelligent Systems (IMPRS-IS) for supporting Jenny Schmalfuss and Victor Oei.
[…]
Supplementary Material
Appendix A Repository Links and Checkpoints
To support reproducibility, Tab. A1 reports the repositories and checkpoints for the optical flow, scene flow and stereo models, which were benchmarked on RobustSpring in Tab. 2, Tab. 4 and Tab. 3. Further details on training and checkpoints for these models can be found in their original publications.

### v2

@@ 4 Results
We evaluate RobustSpring under two aspects: First, we report initial results for 17 optical flow, scene flow and stereo models. Then, we analyze the benchmark evaluation, particularly subsampling strategy and ranking methods.
@@ 4.1 Initial RobustSpring Benchmark Results
We provide initial results on RobustSpring for selected models from all three dense matching tasks. For optical flow, we include SEA-RAFT (Wang et al., 2024b), GMFlow (Xu et al., 2022b), MS-RAFT+ (Jahedi et al., 2024), FlowFormer (Huang et al., 2022), GMA (Jiang et al., 2021), SPyNet (Ranjan and Black, 2017), RAFT (Teed and Deng, 2020), FlowNet2 (Ilg et al., 2017), and PWCNet (Sun et al., 2018). For scene flow, we evaluate M-FUSE (Mehl et al., 2023a) and RAFT-3D (Teed and Deng, 2021). For stereo estimation, we evaluate RAFT-Stereo (Lipson et al., 2021), ACVNet (Xu et al., 2022a), LEAStereo (Cheng et al., 2020), and GANet (Zhang et al., 2019). An overview of all models and used checkpoints is in the Appendix in Tab. 4. Importantly, none of these models are fine-tuned to either Spring or RobustSpring data, to assess the generalization capacity of existing models.
Optical Flow. The evaluation results in Tab. 1 and Fig. 4(a) show large robustness variations across corruption types. Weather corruptions, especially rain and snow, degrade performance most, while color corruptions have little effect. Model rankings also vary: FlowNet2 performs poorly overall but is the most resilient to noise (Fig. 4(b)). SEA-RAFT and GMFlow achieve the lowest average $R^{c}_{\text{EPE}}$, while GMA yields the lowest median, as detailed in Sec. 4.2.
@@
(c) Accuracy vs. Robustnes.
Figure 4: Accumulated corruption robustness $R^{c}_{\text{EPE}}$ for optical flow models over all corruptions [left], only noise corruptions [middle], and accuracy vs. robustness [right] with a dashed line representing a quadratic polynomial fit. Small values are robust (and accurate) models. All other corruption classes color (purple), blur (blue), noise (cyan), quality (green), and weather (yellow) are in App. A.4 in Fig. 7, and Fig. 8 shows accuracy vs. Median $R^{c}_{\text{EPE}}$.
To examine accuracy–robustness relations, we compare both in Fig. 4(c). Accurate models tend to be more robust, though no method excels in both dimensions, and the trend is weak and nonlinear (see also median robustness in Fig. 8). Fitting a quadratic regression to Fig. 4(c) highlights this nonlinearity. Accuracy and robustness coincide for some corruption families but diverge substantially for others. This stands in contrast to adversarial robustness, where a strong accuracy–robustness tradeoff is observed (Schmalfuss et al., 2022b). The key to understanding these differences is to consider different corruption types individually. Because accurate models need to process fine details in order to achieve highly accurate predictions, this attention to details can be exploited by noise-based adversarial attacks in Schmalfuss et al. (2022b). For RobustSpring’s non-adversarial corruptions, accurate models are also relatively sensitive to noise corruptions (cf. Fig. 4(b)). However, the average corruption robustness is mostly dominated by weather corruptions (cf. Fig. 4(a)), where more accurate models achieve better robustness values. We identify the reason for this in Sec. 4.2: accurate models retain corruption-induced motion errors around the particles rather than spreading them onto the background (cf. Fig. 5(a)). This aligns with the results in Schmalfuss et al. (2023), which found accurate models to be more robust towards adversarial weather attacks.
Architectural trends also emerge. Transformer models (GMFlow, FlowFormer) perform best overall but struggle with noise, likely due to global matching. Hierarchical models (e.g. MS-RAFT+) show balanced robustness and may benefit from multi-scale feature processing to cope with quality degradations. Stacked models (e.g. SEA-RAFT, FlowNet2) uniquely resist noise, possibly due to their progressive refinement across layers. Overall, architecture influences robustness to specific corruptions, but no paradigm is universally superior.
Table 2: Initial RobustSpring results on corruption robustness of scene flow and stereo disparity models, using corruption robustness $R^{c}_{\text{1px}}$, $R^{c}_{\text{Abs}}$ and $R^{c}_{\text{Dl}}$ between clean and corrupted predictions. Low values indicate robust models. Corresponding Disparity 1 from scene flow models LEAStereo (s) for M-FUSE, and GANet (s) for RAFT-3D in Tab. 2(b). Stereo disparity models use Stereo (s) and KITTI (k) checkpoints.
(a) Initial scene flow evaluation.
@@
(b) Initial stereo disparity evaluation.
@@
Scene Flow. The results for scene flow are in Tab. 2(a), and include optical flow and target frame disparity predictions for M-FUSE and RAFT-3D. M-FUSE generally produces more robust optical flow across corruptions with a lower average $R^{c}_{\text{EPE}}$ than RAFT-3D. But both methods suffer significant performance losses for severe weather like rain and noise-based corruptions, e.g. impulse noise. Interestingly, their robustness does not improve compared to conventional optical flow models. Noise and weather corruptions remain a challenge for Disparity 2 predictions. Here, RAFT-3D consistently achieves lower robustness scores compared to M-FUSE, but conditions like impulse noise or rain still notably affect disparity predictions. Overall, both models have limited robustness, but temporal consistency may contribute to lower robustness scores under several corruption types.
Stereo. Results of stereo disparity estimations are presented in Tab. 2(b). The effect of the different corruptions on the performance is significant, with noise and weather-based corruptions leading to the largest errors, especially for GANet and LEAStereo. In particular, Gaussian and impulse noise introduce extremely large errors, highlighting the sensitivity of stereo models to pixel-level noise. Blur distortions, especially zoom blur, also have a severe impact on all models. In contrast, color-based distortions generally yield smaller errors. RAFT-Stereo shows stronger resilience across most corruption groups, performing better on color and noise based corruption than other models, but also struggles with noise and severe weather effects such as rain and snow.
Table 3: Evaluations of the metrics used in RobustSpring.
(a) Influence of subsampling. We compare robustness evaluations on the full test data (Full) to evaluations on Spring’s original subsampling (Spring), original subsampling without Hero-frames (Spring*), and our refined corruption subsampling (Ours).
@@
(b) Robustness ranking of optical flow models with ranking strategies Average $R^{c}_{\text{EPE}}$, Median $R^{c}_{\text{EPE}}$, and Schulze to summarize results over corruptions. Please note that Schulze does not produce numeric values.
@@
4.2 Metrics and Benchmark Capability
Subsampling. We evaluate RobustSpring’s strict data subsampling by comparing to results on the full test set. As shown in Tab. 3(a), our subsampling strategy produces results that are nearly identical to those that include all pixels in the robustness calculation. We observe the largest discrepancy for Spring’s original subsampling, because it includes a handful of full-resolution Hero-frames. If those frames are also subsampled (Spring*), results align with the full dataset. Overall, our stricter subsampling to 0.05% of all data is not only data efficient but also exact.
Metric Ranking. To explore how ranking strategies influence the optical-flow robustness order, we contrast our three summarization strategies: Average, Median, and Schulze (cf. App. 5). The rankings in Tab. 3(b) notably differ across strategies. The Average differs most from the other rankings. For example, it ranks GMFlow 2nd, which is only 5th on Median and Schulze, suggesting a good performance across corruptions without excessive outliers but no top performance on most corruptions. Interestingly, Median and Schulze rankings are more aligned. As Schulze’s ranking involves complex comparisons of per-corruption rankings and must be globally recomputed for new models, the Median ranking is a cheap approximation to it. The ranking strategy has significant implications for selecting robust models. While SEA-RAFT is optimal across rankings, the rankings accentuate different aspects: overall performance, outlier robustness, or balanced performance in pairwise comparisons. Hence, RobustSpring reports them all.
Robustness on Object Corruptions.
@@ 4.2 Metrics and Benchmark Capability
(a) Stability of corruption robustness $R^{c}_{\text{EPE}}$ on rain corruption. Robustness scores and rankings remain stable even if no rain pixels are in the $R^{c}_{\text{EPE}}$ calculation.
@@
(b) Relative robustness to noise on RobustSpring transfers to noisy real-world KITTI data (Menze and Geiger, 2015) for most optical flow models.
Figure 5: Additional evaluations of RobustSpring’s benchmark character.
In RobustSpring, robustness is defined as stability of predictions under corrupted inputs: models are robust if their outputs remain consistent for clean and corrupted scenes (Eq. 1). A natural question is whether this metric remains valid when corruptions introduce moving objects, such as rain or snow. To test this, we separate the contributions of background and corruption pixels by excluding pixels of objects like rain drops from the score calculation. Object pixels are detected via the value difference $d$ between clean and corrupted images, and excluded if (1d)(1-d) exceeds a threshold. Threshold 0 excludes none (the vanilla $R^{\text{Rain}}_{\text{EPE}}$), while threshold 100 excludes all. Figure 5(a) shows the robustness score if rain is excluded from the calculation, along with bars indicating the amount [%] of excluded pixels. Remarkably, the robustness score changes by at most $5\%$ even when all rain pixels (about $90\%$ of the image) are discarded. High scores for rain or snow (cf. App. A.4.4) thus result mainly from mispredictions in the periphery of altered pixels, not from motion predictions on altered pixels. As scene-wide effects dominate, our stability-based robustness yields consistent rankings suited for broad robustness evaluations.
@@
Robustness in the Real World. Finally, we investigate if RobustSpring’s corruption robustness transfers to the real world. To this end, we select the noisiest 10% KITTI data, estimating noise as in Immerkaer (1996). These noisy KITTI frames have no clean counterparts to calculate corruption robustness $R^{\text{Noise}}_{\text{EPE}}$. Thus, we approximate $R^{\text{Noise}}_{\text{EPE}}$ via the accuracy difference on noisy and non-noisy KITTI frames. To account for model-specific performance differences on Spring and KITTI, we normalize with the clean dataset performance and show the resulting relative robustness $\tfrac{R^{\text{Noise}}_{\text{EPE}}}{\text{EPE}^{\text{Clean}}}$ in Fig. 5(b). Relatively robust models with low scores on RobustSpring are also robust on KITTI and vice versa. The only outlier, FlowFormer, overperforms on KITTI, potentially due to outstanding memorization capacity and exposure to KITTI during training. Because overall noise resilience on RobustSpring qualitatively transfers to KITTI, RobustSpring supports model selection for real-world settings where corruption robustness cannot be measured.
@@ 5 Conclusion
With RobustSpring we introduce an image corruption dataset and benchmark that evaluates the robustness of optical flow, scene flow, and stereo models. We carefully design 20 different image corruptions and integrate them in time, stereo, and depth for a holistic evaluation of dense matching tasks. Furthermore, we establish a corruption robustness metric using clean and corrupted predictions, and compare ranking strategies to unify model results across all 20 corruptions. RobustSpring’s benchmark further supports data-efficient result uploads to an evaluation server. Our initial evaluation of 17 optical flow, scene flow, and stereo models reveals an overall high sensitivity to corrupted images. As our robustness results translate to real-world performance, systematic corruption benchmarks like RobustSpring are crucial to uncover potential model performance improvements.
@@
Acknowledgements
Lukas Mehl, Jenny Schmalfuss, and Victor Oei acknowledge funding by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – Project-ID 251654672 – TRR 161: Quantitative Methods for Visual Computing (B04, A07). Jenny Schmalfuss and Andres Bruhn acknowledge support by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – Project-ID 533085500 - Robust Optical Flow. Shashank Agnihotri and Margret Keuper acknowledge support by the DFG Research Unit 5336 – Learning to Sense (L2S). Jenny Schmalfuss and Victor Oei acknowledge support from the International Max Planck Research School for Intelligent Systems (IMPRS-IS).
Appendix A Appendix
A.1 Discussion of Robustness Definition
Scope of RobustSpring’s robustness definition. Robustness is a multifaceted concept. In RobustSpring we focus on one specific and mathematically established perspective: robustness as prediction stability under input corruptions. This choice is motivated by its independence from ground-truth definitions of optical flow, and by its practical utility in assessing whether models maintain consistent scene estimates in the presence of visual disturbances. Other notions of robustness, such as those directly tied to ground truth, are equally valid but outside the scope of this benchmark.
Accuracy vs. Robustness. RobustSpring explicitly disentangles accuracy and robustness. Accuracy measures the deviation between model predictions and ground truth, which depends on the adopted definition of optical flow—here, the true 3D motion of visible surfaces projected into the 2D image plane (Horn and Schunck, 1981; Baker et al., 2011). Robustness, by contrast, is defined via Lipschitz continuity (Eq. 1) as the stability of predictions under corrupted inputs, independent of ground truth.
Why Ground-truth-free Robustness. Some works define robustness as the difference between predictions on corrupted images and the corresponding ground truth (Ranjan et al., 2019; Agnihotri et al., 2024c). This approach entangles robustness with accuracy and requires ambiguous judgments about how each corruption changes the motion field (e.g. elastic transforms, rain, snow). By adopting a ground-truth-free definition, RobustSpring avoids this ambiguity and aligns with established robustness literature (Schmalfuss et al., 2022b; Tsipras et al., 2019; Taori et al., 2020; Hein and Andriushchenko, 2017; Pauli et al., 2022).
Perfect Accuracy vs. Perfect Robustness. Under our definitions, a perfectly accurate method may achieve poor robustness scores for corruptions that genuinely alter the motion field (e.g. snow, elastic transform), because its predictions change in accordance with the altered scene. Conversely, a perfectly robust method (constant predictions) is maximally stable but inaccurate. This disentanglement is intentional: it reveals cases where accuracy and robustness align versus diverge. Such divergence highlights model behavior that plain accuracy metrics cannot capture.
Practical Relevance. As demonstrated in Fig. 5(a), current models tend to propagate spurious foreground motions introduced by rain into the background, leading to degraded scene estimates. RobustSpring’s robustness metric captures this instability and thus provides valuable insight into real-world performance. The combination of accuracy and robustness offers practitioners a two-axis evaluation, allowing trade-offs between precision and stability depending on the target application.
A.2 Links and Checkpoints for Evaluated Models
We evaluated the original, author provided optical flow, scene flow and stereo methods on the RobustSpring dataset. Tab. 4 reports the repositories and checkpoints for the optical flow, scene flow and stereo models, which were benchmarked on RobustSpring in Tab. 1, Tab. 2(a), and Tab. 2(b). Further details on training and checkpoints for these models can be found in their original publications.
Table 4: Repositories and checkpoints used for evaluating methods in RobustSpring. *The Mixed checkpoint MS-RAFT+ is pretrained on Chairs and Things and then fine-tuned on a mix of Sintel, Viper, KITTI 2015, Things, and HD1k.